# Stability of $L^{2}-$invariants on stratified spaces

### v2

@@ 4.2. Construction
Let $f:\overline{M}\to\overline{N}$ be smooth, strongly stratum preserving. Consider the Thom form $\tau[\widetilde{f}^{*}(\mathbb{B}^{N})]$ of the disc bundle $\widetilde{f}^{*}(\mathbb{B}^{N})$. Then the Hilsum-Skandalis maps
$\begin{split}&{}^{e}HS(f):C^{\infty}(\widetilde{N},\Lambda^{*}({}^{e}T^{*}N)\otimes\mathscr{E})\to C^{\infty}(\widetilde{M},\Lambda^{*}({}^{e}T^{*}M)\otimes\widetilde{f}^{*}\mathscr{E}),\\ &{}^{e}HS(f)\omega:=\pi_{*}\Bigl{(}\tau\bigl{[}\widetilde{f}^{*}(\mathbb{B}^{N})\bigr{]}\wedge\mathscr{J}\circ{}^{e}(\exp\circ\,\mathbb{B}\widetilde{f})^{*}\omega\Bigr{)}\\ &\qquad\qquad{}^{w}HS(f):C^{\infty}(\widetilde{N},\Lambda^{*}({}^{w}T^{*}N)\otimes\mathscr{E})\to C^{\infty}(\widetilde{M},\Lambda^{*}({}^{w}T^{*}M)\otimes\widetilde{f}^{*}\mathscr{E}),\\ &\qquad\qquad{}^{w}HS(f)\omega:=\pi_{*}\Bigl{(}\tau\bigl{[}\widetilde{f}^{*}(\mathbb{B}^{N})\bigr{]}\wedge\mathscr{J}\circ{}^{w}(\exp\circ\,\mathbb{B}\widetilde{f})^{*}\omega\Bigr{)},\end{split}$ (4.7)
@@
The maps commute with the twisted de Rham differentials $d_{\mathscr{E}}$ and $d_{\widetilde{f}^{*}\mathscr{E}}$ on smooth compactly supported sections. The metrics $g_{\mathscr{E}}$ are induced by the wedge metrics $g_{N},g_{M}$ together with the $\ell^{2}\Gamma$ inner product along the fibers of $\mathscr{E}$.
@@
The maps are well-defined by Proposition 4.9. We now prove (4.8). Note first that $(M,\rho_{M}^{-2}g_{M})$ and $(N,\rho_{N}^{-2}g_{N})$ are of bounded geometry. Hence, boundedness of ${}^{e}HS(f)$ is discussed in detail in [Spe21, Proposition 4.5], who extended [HiSk92] to non-compact Riemannian manifolds of bounded geometry. Note here, that the construction of ${}^{e}HS(f)$ differs from $T_{f}$ in [Spe21, (112), (147)] only by the boundary isomorphism $\mathscr{J}$.
@@
are uniformly bounded. Given $\omega\in C_{c}^{\infty}(N,\Lambda^{k}(T^{*}N)\otimes\mathscr{E})$ with compact support and any fixed degree $k$, we compute
${}^{w}HS(f)\rho_{N}^{k}\omega$ $\equiv{}^{e}HS(f)\rho_{N}^{k}\omega$
$=\pi_{*}\Bigl{(}(\exp\circ\,\mathbb{B}\widetilde{f})^{*}\rho_{N}^{k}\cdot\tau\bigl{[}\widetilde{f}^{*}(\mathbb{B}^{N})\bigr{]}\wedge\mathscr{J}\circ(\exp\circ\,\mathbb{B}\widetilde{f})^{*}\omega\Bigr{)}$
$=\rho^{k}_{M}\cdot\pi_{*}\Bigl{(}\frac{(\exp\circ\,\mathbb{B}\widetilde{f})^{*}\rho^{k}_{N}}{\pi^{*}\rho^{k}_{M}}\cdot\tau\bigl{[}\widetilde{f}^{*}(\mathbb{B}^{N})\bigr{]}\wedge\mathscr{J}\circ(\exp\circ\,\mathbb{B}\widetilde{f})^{*}\omega\Bigr{)}.$
The presence of the bounded factor $\frac{(\exp\circ\,\mathbb{B}\widetilde{f})^{*}\rho^{k}_{N}}{\pi^{*}\rho^{k}_{M}}$ does not affect the argument in [Spe21, Proposition 4.5] and hence $\rho^{-k}_{M}\cdot{}^{w}HS(f)\rho_{N}^{k}$ is again bounded on the $L^{2}$ completions of $k$-th degree twisted differential forms with respect to $\rho_{N}^{-2}g_{\mathscr{E}}$ and $\rho_{M}^{-2}g_{\mathscr{E}}$. This is equivalent to boundedness of ${}^{w}HS(f)$ in (4.8) as claimed. The last claim that ${}^{e/w}HS(f)$ commutes with the de Rham differentials is proved by [ALMP12, Proposition 9.3 (i)], where Lemma 4.10 is implicitly used. ∎
@@ 4.3. Hilsum-Skandalis maps for homotopic maps
Consider a continuous family $f_{s}:\overline{M}\to\overline{N}$ of smooth strongly stratum preserving maps. Consider the pullback bundles $\widetilde{f}_{s}^{*}(\mathbb{B}^{N})$. Homotopic maps induce vector bundle isomorphisms, and hence there exists a continuous family $A_{s}:\widetilde{f}_{0}^{*}(\mathbb{B}^{N})\to\widetilde{f}_{s}^{*}(\mathbb{B}^{N})$ of vector bundle isomorphisms, leading to the following commutative diagram.
{tikzpicture}
@@ 5. Stability of $L^{2}$-Betti numbers and Novikov-Shubin invariants
$(\mathscr{D}_{\min/\max}(M,f^{*}\mathscr{E}_{\overline{N}}),d_{f^{*}\mathscr{E}_{\overline{N}}})\ \textup{and}\ (\mathscr{D}_{\min/\max}(N,\mathscr{E}_{\overline{N}}),d_{\mathscr{E}_{\overline{N}}}).$
@@
Consider the Hilsum-Skandalis replacements ${}^{w}HS(u_{s})^{\prime}$ and ${}^{w}HS(v_{s})^{\prime}$ of the homotopies. By Theorem 4.11 these maps are bounded on the $L^{2}$ completions, commute with the differential by [ALMP12, Proposition 9.3 (i)] and hence define bounded maps between the minimal and maximal Hilbert complexes
${}^{w}HS(u_{s})^{\prime}:(\mathscr{D}_{\min/\max}(N,\mathscr{E}),d_{\mathscr{E}})\to(\mathscr{D}_{\min/\max}(N,\mathscr{E}),d_{\mathscr{E}}),$
${}^{w}HS(v_{s})^{\prime}:(\mathscr{D}_{\min/\max}(M,\mathscr{E}),d_{\mathscr{E}})\to(\mathscr{D}_{\min/\max}(M,\mathscr{E}),d_{\mathscr{E}}).$
@@
$\begin{split}&K_{u}:\mathscr{D}^{*}_{\min/\max}(N,\mathscr{E})\to\mathscr{D}^{*-1}_{\min/\max}(N,\mathscr{E}),\\ &K_{v}:\mathscr{D}^{*}_{\min/\max}(M,\mathscr{E})\to\mathscr{D}^{*-1}_{\min/\max}(M,\mathscr{E}).\end{split}$ (5.2)
If we had ${}^{w}HS(\textup{id}_{\overline{N}})^{\prime}=\textup{Id}$ and ${}^{w}HS(\textup{id}_{\overline{M}})^{\prime}=\textup{Id}$, this would already mean that the $\mathscr{N}\Gamma$-Hilbert complexes $(\mathscr{D}_{\min/\max}(M,\mathscr{E}),d_{\mathscr{E}})$ and $(\mathscr{D}_{\min/\max}(N),d_{\mathscr{E}})$ are chain homotopy equivalent as in Definition 5.1. This would finish the proof. However, this is not the case. In order to compute $({}^{w}HS(\textup{id}_{\overline{N}})^{\prime}-\textup{Id})$ as well as $({}^{w}HS(\textup{id}_{\overline{M}})^{\prime}-\textup{Id})$, we consider the exponential maps as above (recall that the disc fibres in the bundles $\mathbb{B}^{N}$ and $\mathbb{B}^{M}$ are of radius $\delta$, some lower bound for the injectivity radius)
@@
$\begin{split}&{}^{w}HS(\textup{id}_{\overline{N}})^{\prime}-\textup{Id}=d_{\mathscr{E}}K^{\prime}_{u}+K^{\prime}_{u}d_{\mathscr{E}},\qquad K^{\prime}_{u}:\mathscr{D}^{*}_{\min/\max}(N,\mathscr{E})\to\mathscr{D}^{*-1}_{\min/\max}(N,\mathscr{E}),\\ &{}^{w}HS(\textup{id}_{\overline{M}})^{\prime}-\textup{Id}=d_{\mathscr{E}}K^{\prime}_{v}+K^{\prime}_{v}d_{\mathscr{E}}\qquad K^{\prime}_{v}:\mathscr{D}^{*}_{\min/\max}(M,\mathscr{E})\to\mathscr{D}^{*-1}_{\min/\max}(M,\mathscr{E}).\end{split}$ (5.3)
@@ 5.1. Proof of stability
$b^{*}_{(2),\mathrm{max}/\mathrm{min}}(\overline{N}_{\Gamma})$ $=b^{*}_{(2),\mathrm{max}/\mathrm{min}}(\overline{N},\mathcal{E}_{\overline{N}_{\Gamma}})=b^{*}_{(2),\mathrm{max}/\mathrm{min}}(\overline{M},f^{*}\mathcal{E}_{\overline{N}_{\Gamma}}),$
$\alpha_{*,\mathrm{max}/\mathrm{min}}(\overline{N}_{\Gamma})$ $=\alpha_{*,\mathrm{max}/\mathrm{min}}(\overline{N},\mathcal{E}_{\overline{N}_{\Gamma}})=\alpha_{*,\mathrm{max}/\mathrm{min}}(\overline{M},f^{*}\mathcal{E}_{\overline{N}_{\Gamma}}).$
@@
$b^{*}_{(2),\mathrm{max}/\mathrm{min}}(\overline{M},f^{*}\mathcal{E}_{\overline{N}_{\Gamma}})$ $=b^{*}_{(2),\mathrm{max}/\mathrm{min}}(\overline{M},\mathcal{E}_{f^{*}\overline{N}_{\Gamma}}),$
$\alpha_{*,\mathrm{max}/\mathrm{min}}(\overline{M},f^{*}\mathcal{E}_{\overline{N}_{\Gamma}})$ $=\alpha^{*}_{*,\mathrm{max}/\mathrm{min}}(\overline{M},\mathcal{E}_{f^{*}\overline{N}_{\Gamma}}).$
{tikzpicture}
@@
$b^{*}_{(2),\mathrm{max}/\mathrm{min}}(\overline{N},\mathcal{E}_{f^{*}\overline{N}_{\Gamma}})=b_{(2),\mathrm{max}/\mathrm{min}}(\overline{M},\mathcal{E}_{\overline{M}_{\Gamma}})$ $=b^{*}_{(2),\mathrm{max}/\mathrm{min}}(\overline{M}_{\Gamma}),$
$\alpha_{*,\mathrm{max}/\mathrm{min}}(\overline{N},\mathcal{E}_{f^{*}\overline{N}_{\Gamma}})=\alpha_{*,\mathrm{max}/\mathrm{min}}(\overline{M},\mathcal{E}_{\overline{M}_{\Gamma}})$ $=\alpha_{*,\mathrm{max}/\mathrm{min}}(\overline{M}_{\Gamma}),$
since $\overline{M}_{\Gamma}$ and $f^{*}\overline{N}_{\Gamma}$ are isomorphic as Galois $\Gamma$-coverings. ∎

### v3

@@ 4.2. Construction
Let $f:\overline{M}\to\overline{N}$ be smooth, strongly stratum preserving. Consider thecThom form $\tau[{}^{e}TN]$. Then the Hilsum-Skandalis maps
$\begin{split}&{}^{e}HS(f):C^{\infty}(\widetilde{N},\Lambda^{*}({}^{e}T^{*}N)\otimes\mathscr{E})\to C^{\infty}(\widetilde{M},\Lambda^{*}({}^{e}T^{*}M)\otimes\widetilde{f}^{*}\mathscr{E}),\\ &{}^{e}HS(f)\omega:=\pi_{*}\Bigl{(}(\mathbb{B}\widetilde{f})^{*}\tau\bigl{[}{}^{e}TN\bigr{]}\wedge\mathscr{J}\circ{}^{e}(\exp\circ\,\mathbb{B}\widetilde{f})^{*}\omega\Bigr{)}\\ &\qquad\qquad{}^{w}HS(f):C^{\infty}(\widetilde{N},\Lambda^{*}({}^{w}T^{*}N)\otimes\mathscr{E})\to C^{\infty}(\widetilde{M},\Lambda^{*}({}^{w}T^{*}M)\otimes\widetilde{f}^{*}\mathscr{E}),\\ &\qquad\qquad{}^{w}HS(f)\omega:=\pi_{*}\Bigl{(}(\mathbb{B}\widetilde{f})^{*}\tau\bigl{[}{}^{e}TN\bigr{]}\wedge\mathscr{J}\circ{}^{w}(\exp\circ\,\mathbb{B}\widetilde{f})^{*}\omega\Bigr{)},\end{split}$ (4.10)
@@
The maps commute with the twisted de Rham differentials $d_{\mathscr{E}}$ and $d_{\widetilde{f}^{*}\mathscr{E}}$ on smooth sections. The metrics $g_{\mathscr{E}}$ are induced by the wedge metrics $g_{N},g_{M}$ together with the $\ell^{2}\Gamma$ inner product along the fibers of $\mathscr{E}$.
@@
The maps are well-defined by Proposition 4.9. We now prove (4.11). Note first that $(M,\rho_{M}^{-2}g_{M})$ and $(N,\rho_{N}^{-2}g_{N})$ are of bounded geometry. Hence, boundedness of ${}^{e}HS(f)$ is discussed in detail in [Spe21, Proposition 4.5], who extended [HiSk92] to non-compact Riemannian manifolds of bounded geometry. Note here, that the construction of ${}^{e}HS(f)$ differs from $T_{f}$ in [Spe21, (112), (147)] only by the isomorphism $\mathscr{J}$.
@@
are uniformly bounded. Given now $\omega\in C^{\infty}(\widetilde{N},\Lambda^{k}({}^{w}T^{*}N)\otimes\mathscr{E})$ of any fixed degree $k$, we note that $\rho_{N}^{-k}\omega\in C^{\infty}(N,\Lambda^{k}({}^{e}T^{*}N)\otimes\mathscr{E})$. Therefore, we may compute
${}^{w}HS(f)\omega$ $=\pi_{*}\Bigl{(}(\mathbb{B}\widetilde{f})^{*}\tau\bigl{[}{}^{e}TN\bigr{]}\wedge\mathscr{J}\circ{}^{w}(\exp\circ\,\mathbb{B}\widetilde{f})^{*}(\rho_{N}^{k}\rho_{N}^{-k}\omega)\Bigr{)}$
$=\pi_{*}\Bigl{(}(\mathbb{B}\widetilde{f})^{*}\tau\bigl{[}{}^{e}TN\bigr{]}\wedge\mathscr{J}\circ(\exp\circ\,\mathbb{B}\widetilde{f})^{*}\rho_{N}^{k}\cdot{}^{e}(\exp\circ\,\mathbb{B}\widetilde{f})^{*}(\rho_{N}^{-k}\omega)\Bigr{)}$
$=\rho^{k}_{M}\cdot\pi_{*}\Bigl{(}\alpha\cdot(\mathbb{B}\widetilde{f})^{*}\tau\bigl{[}{}^{e}TN\bigr{]}\wedge\mathscr{J}\circ{}^{e}(\exp\circ\,\mathbb{B}\widetilde{f})^{*}(\rho_{N}^{-k}\omega)\Bigr{)},$
$\textup{where}\ \alpha:=(\exp\circ\,\mathbb{B}\widetilde{f})^{*}\rho^{k}_{N}\cdot\pi^{*}\rho^{-k}_{M}.$
The presence of the bounded factor $\alpha$ does not affect the argument in [Spe21, Proposition 4.5] and hence $\rho^{-k}_{M}\cdot{}^{w}HS(f)\rho_{N}^{k}$ is again bounded on the $L^{2}$ completions of $k$-th degree twisted differential forms with respect to $\rho_{N}^{-2}g_{\mathscr{E}}$ and $\rho_{M}^{-2}g_{\mathscr{E}}$. This is equivalent to boundedness of ${}^{w}HS(f)$ in (4.11) as claimed.
The last claim that ${}^{e/w}HS(f)$ commutes with the de Rham differentials follows, as already observed in [ALMP12, Proposition 9.3 (i)] and [Alb17], from the fact that both maps are compositions of the isometry $\mathscr{J}$, pullback and exterior product by a closed form. The pullbacks commute with the twisted differentials, as noted in Lemma 4.8. The isometry $\mathscr{J}$ commutes with the pullback connections by Lemma 4.10 and hence ${}^{e/w}HS(f)$ commute with the twisted differentials. ∎
@@ 4.3. Hilsum-Skandalis maps for homotopic maps
Consider a continuous family $f_{s}:\overline{M}\to\overline{N}$ of smooth strongly stratum preserving maps. Consider the pullback bundles $\widetilde{f}_{s}^{*}(\mathbb{B}^{N})$. Homotopic maps induce vector bundle isomorphisms, and hence there exists a continuous family $A_{s}:\widetilde{f}_{0}^{*}(\mathbb{B}^{N})\to\widetilde{f}_{s}^{*}(\mathbb{B}^{N})$ of vector bundle isomorphisms, leading to the commutative diagram explained in Fig. 7 below
[…]
@@ 5. Stability of $L^{2}$-Betti numbers and Novikov-Shubin invariants
$(\mathscr{D}_{\max}(M,f^{*}\mathscr{E}_{\overline{N}}),d_{f^{*}\mathscr{E}_{\overline{N}}})\ \textup{and}\ (\mathscr{D}_{\max}(N,\mathscr{E}_{\overline{N}}),d_{\mathscr{E}_{\overline{N}}}).$
@@
Consider the Hilsum-Skandalis replacements ${}^{w}HS(u_{s})^{\prime}$ and ${}^{w}HS(v_{s})^{\prime}$ of the homotopies. By Theorem 4.11 these maps are bounded on the $L^{2}$ completions, commute with the twisted differential on smooth sections and hence act between $\Omega^{*}_{\max}(M,\mathscr{E})$ and $\Omega^{*}_{\max}(N,\mathscr{E})$. Those spaces form core subdomains for $\mathscr{D}^{*}_{\max}(M,\mathscr{E})$and $\mathscr{D}^{*}_{\max}(N,\mathscr{E})$, respectively. Hence we obtain bounded maps between their graph closures, the maximal Hilbert complexes
${}^{w}HS(u_{s})^{\prime}:(\mathscr{D}_{\max}(N,\mathscr{E}),d_{\mathscr{E}})\to(\mathscr{D}_{\max}(N,\mathscr{E}),d_{\mathscr{E}}),$
${}^{w}HS(v_{s})^{\prime}:(\mathscr{D}_{\max}(M,\mathscr{E}),d_{\mathscr{E}})\to(\mathscr{D}_{\max}(M,\mathscr{E}),d_{\mathscr{E}}).$
@@
$\begin{split}&K_{u}:\mathscr{D}^{*}_{\max}(N,\mathscr{E})\to\mathscr{D}^{*-1}_{\max}(N,\mathscr{E}),\\ &K_{v}:\mathscr{D}^{*}_{\max}(M,\mathscr{E})\to\mathscr{D}^{*-1}_{\max}(M,\mathscr{E}).\end{split}$ (5.2)
Indeed we know that $K_{u}:L^{2}\Omega^{*}(N,\mathscr{E})\rightarrow L^{2}\Omega^{*}(N,\mathscr{E})$ is bounded and, by (5.1), we have
$d_{N}K_{u}={}^{w}HS(u_{1})^{\prime}-{}^{w}HS(u_{0})^{\prime}-K_{u}d_{N}.$
Note that the above equality implies immediately that
$K_{u}(\Omega^{*}_{\max}(N,\mathscr{E}))\subset\Omega^{*}_{\max}(N,\mathscr{E})).$
Since, in addition, $K_{u}:L^{2}\Omega^{*}(N,\mathscr{E})\rightarrow L^{2}\Omega^{*}(N,\mathscr{E})$ is bounded and $\Omega^{*}_{\max}(N,\mathscr{E})$ is a core subdomain for $\mathscr{D}^{*}_{\max}(N,\mathscr{E})$, we can conclude that $K_{u}$ preserves $\mathscr{D}^{*}_{\max}(N,\mathscr{E})$. Note that we can not argue in the same way for the minimal domain, since $K_{u}$ does not preserve compact support.
If we had ${}^{w}HS(\textup{id}_{\overline{N}})^{\prime}=\textup{Id}$ and ${}^{w}HS(\textup{id}_{\overline{M}})^{\prime}=\textup{Id}$, this would already mean that the $\mathscr{N}\Gamma$-Hilbert complexes $(\mathscr{D}_{\max}(M,\mathscr{E}),d_{\mathscr{E}})$ and $(\mathscr{D}_{\max}(N),d_{\mathscr{E}})$ are chain homotopy equivalent as in Definition 5.1. This would finish the proof. However, this is not the case. In order to compute $({}^{w}HS(\textup{id}_{\overline{N}})^{\prime}-\textup{Id})$ as well as $({}^{w}HS(\textup{id}_{\overline{M}})^{\prime}-\textup{Id})$, we consider the exponential maps as above (recall that the disc fibres in the bundles $\mathbb{B}^{N}$ and $\mathbb{B}^{M}$ are of radius $\delta$, some lower bound for the injectivity radius)
@@
$\begin{split}&{}^{w}HS(\textup{id}_{\overline{N}})^{\prime}-\textup{Id}=d_{\mathscr{E}}K^{\prime}_{u}+K^{\prime}_{u}d_{\mathscr{E}},\qquad K^{\prime}_{u}:\mathscr{D}^{*}_{\max}(N,\mathscr{E})\to\mathscr{D}^{*-1}_{\max}(N,\mathscr{E}),\\ &{}^{w}HS(\textup{id}_{\overline{M}})^{\prime}-\textup{Id}=d_{\mathscr{E}}K^{\prime}_{v}+K^{\prime}_{v}d_{\mathscr{E}}\qquad K^{\prime}_{v}:\mathscr{D}^{*}_{\max}(M,\mathscr{E})\to\mathscr{D}^{*-1}_{\max}(M,\mathscr{E}).\end{split}$ (5.3)
@@ 5.1. Proof of stability
$b^{*}_{(2),\mathrm{max}}(\overline{N}_{\Gamma})$ $=b^{*}_{(2),\mathrm{max}}(\overline{N},\mathcal{E}_{\overline{N}_{\Gamma}})=b^{*}_{(2),\mathrm{max}}(\overline{M},f^{*}\mathcal{E}_{\overline{N}_{\Gamma}}),$
$\alpha_{*,\mathrm{max}}(\overline{N}_{\Gamma})$ $=\alpha_{*,\mathrm{max}}(\overline{N},\mathcal{E}_{\overline{N}_{\Gamma}})=\alpha_{*,\mathrm{max}}(\overline{M},f^{*}\mathcal{E}_{\overline{N}_{\Gamma}}).$
@@
$b^{*}_{(2),\mathrm{max}}(\overline{M},f^{*}\mathcal{E}_{\overline{N}_{\Gamma}})$ $=b^{*}_{(2),\mathrm{max}}(\overline{M},\mathcal{E}_{f^{*}\overline{N}_{\Gamma}}),$
$\alpha_{*,\mathrm{max}}(\overline{M},f^{*}\mathcal{E}_{\overline{N}_{\Gamma}})$ $=\alpha^{*}_{*,\mathrm{max}}(\overline{M},\mathcal{E}_{f^{*}\overline{N}_{\Gamma}}).$
[…]
@@
$b^{*}_{(2),\mathrm{max}}(\overline{N},\mathcal{E}_{f^{*}\overline{N}_{\Gamma}})=b_{(2),\mathrm{max}}(\overline{M},\mathcal{E}_{\overline{M}_{\Gamma}})$ $=b^{*}_{(2),\mathrm{max}}(\overline{M}_{\Gamma}),$
$\alpha_{*,\mathrm{max}}(\overline{N},\mathcal{E}_{f^{*}\overline{N}_{\Gamma}})=\alpha_{*,\mathrm{max}}(\overline{M},\mathcal{E}_{\overline{M}_{\Gamma}})$ $=\alpha_{*,\mathrm{max}}(\overline{M}_{\Gamma}),$
since $\overline{M}_{\Gamma}$ and $f^{*}\overline{N}_{\Gamma}$ are isomorphic as Galois $\Gamma$-coverings.
Step 4) Since both $\overline{M}$ and $\overline{N}$ are oriented, we have
$*\Delta_{k,\mathrm{rel}}=\Delta_{m-k,\mathrm{abs}}*$
where * denotes the Hodge star operator, $m:=\dim(\overline{M})=\dim(\overline{N})$ and $k\in\{0,...,m\}$. Since
$*:L^{2}\Omega^{k}(\overline{N},\mathcal{E}_{\overline{N}_{\Gamma}})\rightarrow L^{2}\Omega^{m-k}(\overline{N},\mathcal{E}_{\overline{N}_{\Gamma}}),$
$*:L^{2}\Omega^{k}(\overline{M},\mathcal{E}_{\overline{M}_{\Gamma}})\rightarrow L^{2}\Omega^{m-k}(\overline{M},\mathcal{E}_{\overline{M}_{\Gamma}}),$
are isometries of Hilbert spaces we can conclude that
$b^{k}_{(2),\min}(\overline{M}_{\Gamma})=b^{m-k}_{(2),\max}(\overline{M}_{\Gamma})=b^{m-k}_{(2),\max}(\overline{N}_{\Gamma})=b^{k}_{(2),\min}(\overline{N}_{\Gamma})$
$\alpha_{k,\min}(\overline{M}_{\Gamma})=\alpha_{m-k,\max}(\overline{M}_{\Gamma})=\alpha_{m-k,\max}(\overline{N}_{\Gamma})=\alpha_{k,\min}(\overline{N}_{\Gamma}).$